OpenStudy (anonymous):

Give the slope and y-intercept for the equation below. 8x + 2y = 4 m = y-intercept

OpenStudy (experimentx):

first change it to slope intercept form 2y = -8x + 4 y = -8x/2 + 4/2 and you have y = m x + c, m=-8/2, and c=4/2